OpenStudy (anonymous):

How do i convert Cos^5 x into (1/16) [Cos 5 x + 5 cos 3x + 10 cos x] using Euler's Formula

OpenStudy (anonymous):

In regards to Euler I know I take the \[\cos^5 x = (\cos 5x + i \sin 5x) = (\cos x + i \sin x)^5 = (e^{ix})^5\]